Your company is evaluating a purchase of Skeksis Crystal Mining LLC. Skekis produced free cash flow of $7.3m last year, and this is expected to grow at 11% for the next five years before leveling off to a long term growth rate of 2%. Your company has a cost of capital of 10%, while Skeksis has a cost of capital of 13%. Skeksis has 3 million shares outstanding and $25m in debt. What would be the highest price you would pay per share of Skeksis’ stock? Please show ALL work don't compute on excel please!
Answer #1
WACC= 10.00%
Year Previous year FCF FCF growth rate FCF current year Horizon value Total Value Discount factor Discounted value
1 7.3 11.00% 8.103 8.103 1.1 7.3664
2 8.103 11.00% 8.99433 8.99433 1.21 7.43333
3 8.99433 11.00% 9.9837063 9.9837063 1.331 7.50091
4 9.9837063 11.00% 11.08191399 11.08191399 1.4641 7.5691
5 11.08191399 11.00% 12.30092453 156.837 169.1379245 1.61051 105.02134
Long term growth rate (given)= 2.00% Value of Enterprise = Sum of discounted value = 134.89

Alternatively: value of enterprise = 7.3*(1+0.11)/(1+0.1)+7.3*((1+0.11)/(1+0.1))^2+7.3*((1+0.11)/(1+0.1))^3+7.3*((1+0.11)/(1+0.1))^4+7.3*((1+0.11)/(1+0.1))^5+7.3*(1+0.11)^5*(1+0.02)/(0.1-0.02)/(1+0.1)^5 = 134.89

Where
Current FCF =Previous year FCF*(1+growth rate)^corresponding year
Total value = FCF + horizon value (only for last year)
Horizon value = FCF current year 5 *(1+long term growth rate)/( WACC-long term growth rate)
Discount factor=(1+ WACC)^corresponding period
Discounted value=total value/discount factor
Enterprise value = Equity value+ MV of debt
134.89 = Equity value+25
Equity value = 109.89
share price = equity value/number of shares
share price = 109.89/3
share price = 36.63